OpenStudy (anonymous):

Combing Like Terms 2(2x - 4) + x = 7

OpenStudy (anonymous):

4x-8 + x = 7 4x-x= 7 + 8 3x = 15 x = 5

OpenStudy (anonymous):

2(2x5-4)+5=7

OpenStudy (anonymous):

is the 5 near the 2x an exponent

OpenStudy (anonymous):

2(2x^5-4)+5=7 4^6 -8 + 5 =7 4^6 -3 = 7 4^6 = 7 + 3 4^6= 10 i think i did it but idk how to finish